\usepackage[light,first]{draftcopy}

Test2: Should print ENTWURF on the first page, light. This is a test page 1

Test2: Should print ENTWURF on the first page, light. This is a test page 2

This is a testpage 1 This is a testpage 2 This is a testpage 3 This is a testpage 4